OpenStudy (anonymous):

solve by factoring: 4x^2=-4x-1

geerky42 (geerky42):

Add both sides by 4x+1.

OpenStudy (anonymous):

okay so then you get 4x^2+4x+1=0 how do you factor this

OpenStudy (anonymous):

@geerky42

OpenStudy (mathstudent55):

To factor 4x^2 + 4x + 1, set up two sets of parentheses: ( )( ) On the left side of each you need two factors of 4x^2. Either 4x and x, or 2x and 2x. Let's try 2x and 2x (2x )(2x )

OpenStudy (mathstudent55):

On the right side of the parentheses, you need two factors of 1 which have to be 1 and 1: (2x + 1)(2x + 1) Now FOIL it to make sure it's right.

OpenStudy (anonymous):

It is, thank youuuuuuuu, so that makes the answer 1/2 right

OpenStudy (mathstudent55):

Be careful. Write it out. (2x + 1)(2x + 1) = 0 2x + 1 = 0 or 2x + 1 = 0 Subtract 1 from both sides. What do you get?

OpenStudy (anonymous):

i got -1/2

OpenStudy (mathstudent55):

Better.

OpenStudy (anonymous):

oh 2x = -1

OpenStudy (anonymous):

okay thank you

OpenStudy (mathstudent55):

right, so when you now divide both sides by 2 you get x = -1/2